An element is first connected across an external resistance R1=2 Ω, and

then across an external resistance R2= 0.5 Ω. Find the e.m.f. of the

element and its internal resistance if in each of these cases the power

evolved in the external circuit is the same and equal to 2.54 W.


1
Expert's answer
2021-06-21T11:37:54-0400

"I= \frac{E}{R+r}"


"P=I^2R= \frac{E^2R}{(R+r)^2}"


"P_1=\frac{E^2*2}{(2+r)^2}"


"P_2=\frac{E^2*0.5}{(0.5+r)^2}"

"P_1 = P_2=2.54"

"\frac{E^2*2}{(2+r)^2}=\frac{E^2*0.5}{(0.5+r)^2}"

"4*(0.5+r)^2=(2+r)^2"

"2*(0.5+r)=(2+r)"

"r=1"

"P_1=I_1^2R_1"

"I_1=\sqrt{\frac{P_1}{R_1}}=\sqrt{\frac{2.54}{2}}=1.17"

"I_1= \frac{E}{R_1+r}"

"E=I_1(R_1+r)=1.17(2+1)=3.51"

"\text{Answer: }e.m.f = 3.51V;r=1\Omega"